OpenStudy (anonymous):

I need help finding the vertex and axis of symmetry to y=-3x^2

OpenStudy (anonymous):

vertex-0,0 axis of symmetry-0

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

graph it ( i used a graphing calculator) the point where it turns is the origin the y axis splits it evenly in half

OpenStudy (anonymous):

Thank you. I will probably need more help. I hate College INtermediate Algebra.

OpenStudy (anonymous):

Do you want a formula to get the vertex and the axis of symmetry without graphing?

OpenStudy (anonymous):

Anyway, here's the formula. Vertex is (h,k) The standard form in ax^2+ bx+ c= y or 0 Your equation is y= -3x^2. That makes a= -3, b=0 and c=0 as well. Formula for h is -b/2a Substitute and you have h= -0/2(-3) h= 0 Formula for k is c-h^2a Substitute and you have k= 0-0^2(-3) k=0 Vertex now is (0,0) Now the axis of symmetry is equal to x. And x is equal to h. Substitute and you have x=0. Lol, there. Without graphing. k= 0